OpenStudy (anonymous):

An airplane ascends at the rate of 40 feet per minute. The function below shows the height of the airplane from the ground f(t) after t minutes. f(t) = 40t + 520 Which statement best describes the y intercept of the graph of the function?

OpenStudy (anonymous):

It is equal to 520 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 560 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 40 and represents the height of the airplane above the ground, in feet, before it starts ascending. It is equal to 520 and represents the height of the airplane above the ground, in feet, before it starts ascending.

OpenStudy (anonymous):

I think the answer is D.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I know the answer is either A or B.

OpenStudy (phi):

and the y intercept is when x=0 or in this case t=0

OpenStudy (anonymous):

Yeah which means: f(x) = 40t + 520 f(x) = 40(0) + 520 f(x) = 0 + 520 f(x) = 520 y = 520

OpenStudy (anonymous):

Yeah but which explanation? I think D's explanation makes more sense.

OpenStudy (phi):

in A, "after 1 minute" means t=1 but t = 0

OpenStudy (anonymous):

Oh, okay.

OpenStudy (phi):

in D, "before it starts ascending" means the plane stays at 520 until it reaches time 0 when it starts going up. D is the answer